You are here

Beam delivery update: Beam schedule changes

header

04 December 2014

As was the case last night, the cyclotron’s delivering ~125 µA to BL1A, 9.8 µA to BL2A, and nA as needed to BL2C1 for PIF. This is expected to continue overnight and through tomorrow morning. BL2A will tentatively go off for two hours around 13:00 to allow access to the ISAC mass separator room, but once 2A resumes operation we expect to run as we are now though the weekend.

At ISAC, 46K was delivered to GRIFFIN overnight last night; the experiment switched to 47K at about 09:00 this morning. GRIFFIN will continue to take beam through the weekend. We’ve confirmed that we’ll carry out a 213Fr implantation for life sciences experiment L112 on Monday, with an option to do that Tuesday night if needed; the online beam schedule has been updated to reflect this. The ISAC RIB schedule hasn’t yet been changed beyond that: we plan to deliver beam for francium trapping as scheduled, and a decision regarding the subsequent polarizer run will be made once we’re able to measure Mg yields during the setup for L112 on Monday.

Also at ISAC, the OLIS multi-charge source will be restarted by ISAC Ops tonight in preparation for the delivery of 76Se to DRAGON tomorrow. This will be used only to measure charge-state distributions as problems with the first electric dipole in DRAGON preclude taking data for the experiment that was planned. The beam schedule has been updated to reflect this as well.

I should note that a couple of other housekeeping-type changes were made to the schedule this afternoon: Tuesday’s maintenance shift has been reclassified as beam development, and isotope production at STF has been scheduled for December 9–16.

-- ACM